# WAFT: Warping-Alone Field Transforms for Optical Flow

### v1

@@ Abstract
We introduce Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy with lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance. WAFT is a simple and flexible meta-architecture with minimal inductive biases and reliance on custom designs. Compared with existing methods, WAFT ranks 1st on Spring and KITTI benchmarks, achieves the best zero-shot generalization on KITTI, while being up to 4.1×4.1$\times$4.1 × faster than methods with similar performance. Code and model weights are available at https://github.com/princeton-vl/WAFT.
@@ 1 Introduction
Optical flow is a fundamental level vision task that estimates per-pixel 2D motion between video frames. It has many downstream applications, including 3D reconstruction and synthesis [29, 68], action recognition [49, 38, 65], frame interpolation [59, 25, 14], and autonomous driving [9, 33, 19].
Cost volumes [46, 16] with iterative updates [50, 52] has become a standard design in most state-of-the-art methods [46, 8, 58, 50, 13, 52, 34], especially when both accuracy and efficiency are taken into account. Previous work [46, 50] regards cost volume as a more effective representation than image features, as it explicitly models the visual similarity between pixels.
However, constructing cost volumes is expensive in both time and memory [64, 55]. The cost increases quadratically with the radius of the neighborhood. As a result, cost volumes are often constructed from low resolution features, limiting the ability of the model to handle high-resolution input images.
@@
The design of WAFT is simple, with flow-specific designs kept to the minimum. WAFT consists of an input encoder that extracts features from individual input frames and a recurrent update module that iteratively updates flow. Compared to other RAFT-like architectures, WAFT is much simplified because it does not use cost volumes and has removed the context encoder that provides extra features for the update module. WAFT is designed to function as a meta-architecture for optical flow in the sense that the individual components including input encoder and the update unit do not require custom designs and can use existing off-the-shelf (pretrained) architectures. In our experiments, we evaluate different choices (such as ResNet [11] and DPT [39]) that have different accuracy-efficiency trade-offs.
WAFT achieves state-of-the-art performance across various benchmarks with high efficiency and a simple design. On KITTI [9], WAFT achieves the best Fl on non-occluded pixels and the second best on all pixels. On Spring [31], WAFT ranks first across all metrics. It also achieves the best zero-shot cross-dataset generalization on KITTI. We achieve this with standard network architectures for the sub-modules  [7, 11, 39], removing custom designs typically needed in prior work, while being up to 4.1$\times$ faster than existing methods.
Our main contributions are two fold: (1) we challenge the conventional wisdom that cost volume is a key component for achieving state-of-the-art accuracy and efficiency for optical flow; (2) we introduce WAFT, a warping-based meta-architecture that is simpler and achieves state-of-the-art accuracy with high efficiency.
Figure 1: The meta-architecture of WAFT consists of an input encoder and a recurrent update module. We first extract image features from the input encoder, and then use these features to iteratively update the flow estimate for $T$ steps. At each step, we perform feature indexing through a lightweight backward warping on the feature of frame 2, removing the dependency on expensive cost volume used by previous work [46, 50].
@@ 2 Related Work
Traditional methods treated optical flow as a global optimization problem that maximizes visual similarity between corresponding pixels [12, 61, 4, 2]. These methods apply coarse-to-fine warping [2, 1, 32], a strategy theoretically justified by Brox et al. [2], to solve this optimization.
Today, this field is dominated by deep learning methods [16, 8, 46, 62, 15, 50, 45, 48, 5, 13, 44, 53, 54, 56, 57, 24, 43, 17, 18, 27, 66, 63, 28, 21, 26, 67, 34], which can be categorized into two paradigms: direct or iterative. Direct methods [8, 53, 54, 43, 24, 56] treat flow estimation as a standard dense prediction task (e.g. monocular depth estimation) and directly regress the dense flow field from large-scale pre-trained models. Iterative methods [50, 52, 26, 34, 67, 46, 13] align more closely with traditional warping-based approaches, refining the flow predictions progressively. Most state-of-the-art methods [50, 52, 13, 44, 26, 34] follow the iterative paradigm due to its significantly higher efficiency than the direct ones.
Cost volumes [46, 8] have been regarded as a standard design in iterative methods [50, 52, 34, 13, 44, 57, 26]. Prior work [50, 46] empirically shows the effectiveness of cost volumes in handling large displacements. Many iterative methods [50, 52, 34] adopt partial cost volumes [46] to avoid the quadratic computational complexity of full 4D cost volumes; they restrict the search range of each pixel in frame 1 to the neighborhood of its corresponding pixel in frame 2. However, they still suffer from the high memory consumption inherent in cost volumes [55, 64].
WAFT is a warping-based iterative method. We achieve state-of-the-art performance across various benchmarks without constructing cost volumes, challenging the conventional wisdom established by previous work [46, 50, 13]. Warping no longer suffers from high memory consumption inherent in cost volumes, which enables high-resolution indexing and therefore improves accuracy.
@@
Vision transformers [7] have achieved significant progress across a wide range of visual tasks [60, 22, 37, 10, 42]. In the context of optical flow, most direct methods [53, 54, 43] regress flow from a large-scale pre-trained vision transformer with a lightweight flow head. Many iterative methods [13, 44, 26, 67] design task-specific transformer blocks to process cost volumes.
WAFT adopts similar designs to DPT [39] in its recurrent update module, which implicitly handles large displacements in optical flow through the transformer architecture. We empirically show that this is crucial to make warping work. WAFT can also benefit from large-scale pre-trained transformers like existing methods [43, 53, 54, 67], with minimal additional flow-specific designs.
@@ 3.1 Iterative Methods with Cost Volume
Given two adjacent RGB frames, optical flow predicts pixel-wise 2D motion between adjacent frames. Current iterative methods [50, 13, 52] consist of two parts: (1) input encoders that extract dense image features at low resolution, and (2) a recurrent update module that iteratively refines the flow estimate.
@@
where $\cdot$ represents the dot product of two vectors. At each step, the recurrent update module indexes into the cost volume using the current flow estimate and predicts the residual flow update. Several methods [13, 44] directly process $V$ for better performance, but tend to be more costly.
Partial cost volume [46] is introduced to reduce the cost by avoiding the construction of a full 4D cost volume. Given the current flow estimate $f_{cur}\in\mathbb{R}^{h\times w\times 2}$ and a pre-defined look-up radius $r$, partial cost volume $V_{par}:\mathbb{R}^{h\times w\times 2}\to\mathbb{R}^{h\times w\times(r^{2})}$ implements an on-the-fly partial construction by restricting the indexing range of a pixel $p\in I_{1}$ to the neighborhood of its corresponding pixel $p+(f_{cur})_{p}\in I_{2}$, formulated as:
@@
where the operator “concat” concatenates all values inside the set into a vector. In practice, partial cost volumes are usually constructed at multiple scales [50, 52] to improve the prediction of large displacements. Current methods also introduce context encoder [46, 50, 13, 52] to enhance the effectiveness of iterative refinement.
@@ 3.2 Drawbacks of Cost Volumes
Figure 3: Visualizations of different methods [52, 34] on Spring [31]. WAFT, benefiting from high-resolution indexing, obtains sharper boundaries and lower errors than low-resolution approaches.
@@
SEA-RAFT [52] 14.1 25.8 OOM
WAFT (Ours) - - 9.5
@@
The main drawback of cost volume is its high memory consumption. Full or partial cost volume at high resolution are very expensive and often infeasible [64, 55]. Therefore, most iterative methods build the cost volume and index into it at 1/8 resolution. To further demonstrate the problem, we implement several variants of SEA-RAFT [52] that build partial cost volumes at different resolutions. We set the base channel dimension as 32, 64, and 128 for the 1/2, 1/4, and 1/8 resolution variants, respectively, to make their computational cost similar (around 350GMACs). As shown in Table 1, the partial cost volume with look-up radius $r=4$ in SEA-RAFT runs out of memory at 1/2 resolution.
@@
Since cost volumes are restricted to low resolution, the predicted flow field must be downsampled for cost volume look-up, which inevitably introduces errors. As illustrated in Figure 3 using an example from the Spring benchmark [31], existing methods struggle to produce clear boundaries, particularly noticeable in the top-right corner. In contrast, our method benefits from high-resolution look-up of feature vectors and obtains sharper boundary predictions in these challenging regions. We also show the quantitative performance gain from high-resolution indexing in Table 5.
@@
Warping was widely used in both classical and early deep learning approaches [32, 2, 16, 40]. However, most recent iterative methods [52, 50, 13] have either replaced warping with cost volumes or used both in combination, since cost volumes have been shown to remarkably improve the performance [46]. In this section, we analyze the similarities and differences between warping and cost volume (see Figure 2), and argue that with appropriate designs, warping-based methods can achieve performance on par with cost-volume-based methods.
The overlap of warping and cost volumes lies in their use of the current flow prediction to index into feature maps, which is closely related to optimization [2]. In cost-volume-based iterative methods, the current flow estimate $f_{cur}$ is used to define partial cost volume $V_{par}$ (see Section 3.1). It is also used to define the warped feature map $\texttt{Warp}(f_{\text{cur}})\in\mathbb{R}^{h\times w\times d}$, where the feature vector of pixel $p\in I_{1}$ is indexed from the feature map of frame 2, formulated as:
@@
Compared to cost volumes, for each pixel in frame 1, warping does not calculate its visual similarity to multiple pixels in frame 2, making it no longer able to explicitly model large displacements. However, we can implicitly handle this long-range dependence through the attention mechanism in vision transformers [7, 39], which, as we will demonstrate, is crucial to make warping work well (see Section 5). The high memory efficiency of warping also enables high-resolution indexing, leading to improved accuracy.
@@ 4 Method
Figure 4: Visualizations on Spring [31], KITTI [9], and Sintel [3].
@@
The input encoder $F$ takes a single image $I$ as input and outputs a feature map. The input encoder can be based on current large-scale pre-trained models. We use frozen features from DepthAnythingV2 [60], and further refine them with a ResNet [11]. Specifically, we concatenate the feature from DepthAnythingV2 [60] with the original image and feed it into the ResNet [11], followed by a deconvolution head which upsamples the low-resolution ResNet feature to half of the original image resolution: i.e. $(h,w)=(\frac{H}{2},\frac{W}{2}),F(I)\in\mathbb{R}^{h\times w\times d}$.
@@
Similar to existing iterative methods [50, 52, 13], our recurrent update module $R$ iteratively predicts the residual flow updates. At step $t$, we concatenate $F(I_{1})$ (feature of frame 1), $\texttt{Warp}(f_{\text{cur}})$ (warped feature of frame 2, see Section 3.3), and the current hidden state $\text{Hidden}_{t}\in\mathbb{R}^{h\times w\times d}$ as input. We use a slightly modified DPT [39] as the architecture of the module.
@@
We adopt the Mixture-of-Laplace (MoL) loss used in SEA-RAFT [52]. At step $t$, the hidden state $\text{Hidden}_{t}$ is used to predict the MoL parameters $M\in\mathbb{R}^{h\times w\times 6}$. They are upsampled to the original image resolution through convex upsampling [50, 52].
@@
We replace cost volumes, which are standard in existing iterative methods [46, 50, 13, 52], with high-resolution warping, which is more memory-efficient. In addition, we have removed the context encoder [46], another flow-specific design standard in existing iterative methods.
A direct benefit from our simplified design is that we can load pre-trained weights for standard architectures such as DPT [7, 39], which can improve generalization as our experiments will show Section 5.
The simplicity of our meta-architecture also enables more apples-to-apples comparisons between direct methods and iterative methods. Existing direct methods [53, 54, 43] share an input format similar to that of the first iteration of WAFT, making them more directly compatible. We will empirically show the effectiveness and necessity of iterative indexing within our meta-architecture in Section 5.

### v2

@@ Abstract
We introduce Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy with lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance. WAFT is a simple and flexible meta-architecture with minimal inductive biases and reliance on custom designs. Compared with existing methods, WAFT ranks 1st on Spring, Sintel, and KITTI benchmarks, achieves the best zero-shot generalization on KITTI, while being up to $4.1\times$ faster than methods with similar performance. Code and model weights are available at https://github.com/princeton-vl/WAFT.
@@ 1 Introduction
Optical flow is a fundamental level vision task that estimates per-pixel 2D motion between video frames. It has many downstream applications, including 3D reconstruction and synthesis (Ma et al., 2022; Zuo & Deng, 2022), action recognition (Sun et al., 2018b; Piergiovanni & Ryoo, 2019; Zhao et al., 2020b), frame interpolation (Xu et al., 2019; Liu et al., 2020; Huang et al., 2020), and autonomous driving (Geiger et al., 2013; Menze & Geiger, 2015; Janai et al., 2020).
Cost volumes (Sun et al., 2018a; Ilg et al., 2017) with iterative updates (Teed & Deng, 2020; Wang et al., 2024) has become a standard design in most state-of-the-art methods (Sun et al., 2018a; Dosovitskiy et al., 2015; Xu et al., 2017; Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024; Morimitsu et al., 2025), especially when both accuracy and efficiency are taken into account. Previous work (Sun et al., 2018a; Teed & Deng, 2020) regards cost volumes as a more effective representation than image features, as it explicitly models the visual similarity between pixels.
However, constructing cost volumes is expensive in both time and memory (Zhao et al., 2024; Xu et al., 2023a). The cost increases quadratically with the radius of the neighborhood. As a result, cost volumes are often constructed from low resolution features, limiting the ability of the model to handle high-resolution input images.
@@
The design of WAFT is simple, with flow-specific designs kept to the minimum. WAFT consists of an input encoder that extracts features from individual input frames and a recurrent update module that iteratively updates flow. Compared to other RAFT-like architectures, WAFT is much simplified because it does not use cost volumes and has removed the context encoder that provides extra features for the update module. WAFT is designed to function as a meta-architecture for optical flow in the sense that the individual components including input encoder and the update unit do not require custom designs and can use existing off-the-shelf (pretrained) architectures. In our experiments, we evaluate different choices (such as ResNet (He et al., 2016) and DPT (Ranftl et al., 2021)) that have different accuracy-efficiency trade-offs.
[…]
Our main contributions are two-fold: (1) we challenge the conventional wisdom that cost volume is a key component for achieving state-of-the-art accuracy and efficiency for optical flow; (2) we introduce WAFT, a warping-based meta-architecture that is simpler and achieves state-of-the-art accuracy with high efficiency.
Figure 1: The meta-architecture of WAFT consists of an input encoder and a recurrent update module. We first extract image features from the input encoder, and then use these features to iteratively update the flow estimate for $T$ steps. At each step, we perform feature indexing through a lightweight backward warping on the feature of frame 2, removing the dependency on expensive cost volume used by previous work.
@@ 2 Related Work
Traditional methods treated optical flow as a global optimization problem that maximizes visual similarity between corresponding pixels (Horn & Schunck, 1981; Zach et al., 2007; Chen & Koltun, 2016; Brox et al., 2004). These methods apply coarse-to-fine warping (Brox et al., 2004; Black & Anandan, 1996; Memin & Perez, 1998), a strategy theoretically justified by Brox et al(Brox et al., 2004), to solve this optimization.
Today, this field is dominated by deep learning methods (Ilg et al., 2017; Dosovitskiy et al., 2015; Sun et al., 2018a; Zhao et al., 2020a; Hui et al., 2018; Teed & Deng, 2020; Sui et al., 2022; Sun et al., 2022; Deng et al., 2023; Huang et al., 2022; Shi et al., 2023; Weinzaepfel et al., 2022; 2023; Xu et al., 2022; 2023b; Leroy et al., 2023; Saxena et al., 2024; Jahedi et al., 2024; 2023; Luo et al., 2022; Zheng et al., 2022; Zhao et al., 2022; Luo et al., 2023; Jung et al., 2023; Luo et al., 2024; Zhou et al., 2024; Morimitsu et al., 2025), which can be categorized into two paradigms: direct or iterative. Direct methods (Dosovitskiy et al., 2015; Weinzaepfel et al., 2022; 2023; Saxena et al., 2024; Leroy et al., 2023; Xu et al., 2022) treat flow estimation as a standard dense prediction task (e.g. monocular depth estimation) and directly regress the dense flow field from large-scale pre-trained models. Iterative methods (Teed & Deng, 2020; Wang et al., 2024; Luo et al., 2024; Morimitsu et al., 2025; Zhou et al., 2024; Sun et al., 2018a; Huang et al., 2022) align more closely with traditional warping-based approaches, refining the flow predictions progressively. Most state-of-the-art methods (Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022; Shi et al., 2023; Luo et al., 2024; Morimitsu et al., 2025) follow the iterative paradigm due to its significantly higher efficiency than the direct ones.
Cost volumes (Sun et al., 2018a; Dosovitskiy et al., 2015) have been regarded as a standard design in iterative methods (Teed & Deng, 2020; Wang et al., 2024; Morimitsu et al., 2025; Huang et al., 2022; Shi et al., 2023; Xu et al., 2023b; Luo et al., 2024). Prior work (Teed & Deng, 2020; Sun et al., 2018a) empirically shows the effectiveness of cost volumes in handling large displacements. Many iterative methods (Teed & Deng, 2020; Wang et al., 2024; Morimitsu et al., 2025) adopt partial cost volumes (Sun et al., 2018a) to avoid the quadratic computational complexity of full 4D cost volumes; they restrict the search range of each pixel in frame 1 to the neighborhood of its corresponding pixel in frame 2. However, they still suffer from the high memory consumption inherent in cost volumes (Xu et al., 2023a; Zhao et al., 2024).
WAFT is a warping-based iterative method. We achieve state-of-the-art performance across various benchmarks without constructing cost volumes, challenging the conventional wisdom established by previous work (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022). Warping no longer suffers from high memory consumption inherent in cost volumes, which enables high-resolution indexing and therefore improves accuracy.
@@
Vision transformers (Dosovitskiy et al., 2020) have achieved significant progress across a wide range of visual tasks (Yang et al., 2024; Kirillov et al., 2023; Oquab et al., 2023; He et al., 2022; Rombach et al., 2022). In the context of optical flow, most direct methods (Weinzaepfel et al., 2022; 2023; Saxena et al., 2024) regress flow from a large-scale pre-trained vision transformer with a lightweight flow head. Many iterative methods (Huang et al., 2022; Shi et al., 2023; Luo et al., 2024; Zhou et al., 2024) design task-specific transformer blocks to process cost volumes.
WAFT adopts similar designs to DPT (Ranftl et al., 2021) in its recurrent update module, which implicitly handles large displacements in optical flow through the transformer architecture. We empirically show that this is crucial to make warping work. WAFT can also benefit from large-scale pre-trained transformers like existing methods (Saxena et al., 2024; Weinzaepfel et al., 2022; 2023; Zhou et al., 2024), with minimal additional flow-specific designs.
@@ 3.1 Iterative Methods with Cost Volume
Given two adjacent RGB frames, optical flow predicts pixel-wise 2D motion between adjacent frames. Current iterative methods (Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024) consist of two parts: (1) input encoders that extract dense image features at low resolution, and (2) a recurrent update module that iteratively refines the flow estimate.
@@
where $\cdot$ represents the dot product of two vectors. At each step, the recurrent update module indexes into the cost volume using the current flow estimate and predicts the residual flow update. Several methods (Huang et al., 2022; Shi et al., 2023) directly process $V$ for better performance, but tend to be more costly.
Partial cost volume (Sun et al., 2018a) is introduced to reduce the cost by avoiding the construction of a full 4D cost volume. Given the current flow estimate $f_{cur}\in\mathbb{R}^{h\times w\times 2}$ and a pre-defined look-up radius $r$, partial cost volume $V_{par}:\mathbb{R}^{h\times w\times 2}\to\mathbb{R}^{h\times w\times(r^{2})}$ implements an on-the-fly partial construction by restricting the indexing range of a pixel $p\in I_{1}$ to the neighborhood of its corresponding pixel $p+(f_{cur})_{p}\in I_{2}$, formulated as:
@@
where the operator “concat” concatenates all values inside the set into a vector. In practice, partial cost volumes are usually constructed at multiple scales (Teed & Deng, 2020; Wang et al., 2024) to improve the prediction of large displacements. Current methods also introduce context encoder (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024) to enhance the effectiveness of iterative refinement.
@@ 3.2 Drawbacks of Cost Volumes
Figure 3: Visualizations of different methods on Spring (Mehl et al., 2023). WAFT, benefiting from high-resolution indexing, obtains sharper boundaries and lower errors than low-resolution approaches.
@@
SEA-RAFT 14.1 25.8 OOM
Flowformer 26.1 - -
CCMR+ 36.0 - -
WAFT-Twins-a2 7.0 7.6 9.2
@@
The main drawback of cost volume is its high memory consumption. Full or partial cost volume at high resolution are very expensive and often infeasible (Zhao et al., 2024; Xu et al., 2023a). Therefore, most iterative methods build the cost volume and index into it at 1/8 resolution. To further demonstrate the problem, we implement several variants of SEA-RAFT (Wang et al., 2024) that build partial cost volumes at different resolutions. We set the base channel dimension as 32, 64, and 128 for the 1/2, 1/4, and 1/8 resolution variants, respectively, to make their computational cost similar (around 350GMACs). As shown in Table 1, the partial cost volume with look-up radius $r=4$ in SEA-RAFT runs out of memory at 1/2 resolution. WAFT removes the reliance on cost volume, and therefore consumes significantly lower memory than existing methods (Wang et al., 2024; Huang et al., 2022; Jahedi et al., 2024).
@@
Since cost volumes are restricted to low resolution, the predicted flow field must be downsampled for cost volume look-up, which inevitably introduces errors. As illustrated in Figure 3 using an example from the Spring benchmark (Mehl et al., 2023), existing methods struggle to produce clear boundaries, particularly noticeable in the top-right corner. In contrast, our method benefits from high-resolution look-up of feature vectors and obtains sharper boundary predictions in these challenging regions. We also show the quantitative performance gain from high-resolution indexing in Table 5.
@@
Warping was widely used in both classical and early deep learning approaches (Memin & Perez, 1998; Brox et al., 2004; Ilg et al., 2017; Ranjan & Black, 2017). However, most recent iterative methods (Wang et al., 2024; Teed & Deng, 2020; Huang et al., 2022) have either replaced warping with cost volumes or used both in combination, since cost volumes have been shown to remarkably improve the performance (Sun et al., 2018a). In this section, we analyze the similarities and differences between warping and cost volume (see Figure 2), and argue that with appropriate designs, warping-based methods can achieve performance on par with cost-volume-based methods.
The overlap of warping and cost volumes lies in their use of the current flow prediction to index into feature maps, which is closely related to optimization (Brox et al., 2004). In cost-volume-based iterative methods, the current flow estimate $f_{cur}$ is used to define partial cost volume $V_{par}$ (see Section 3.1). It is also used to define the warped feature map $\texttt{Warp}(f_{\text{cur}})\in\mathbb{R}^{h\times w\times d}$, where the feature vector of pixel $p\in I_{1}$ is indexed from the feature map of frame 2, formulated as:
@@
Compared to cost volumes, for each pixel in frame 1, warping does not calculate its visual similarity to multiple pixels in frame 2, making it no longer able to explicitly model large displacements. However, we can implicitly handle this long-range dependence through the attention mechanism in vision transformers (Dosovitskiy et al., 2020; Ranftl et al., 2021), which, as we will demonstrate, is crucial to make warping work well (see Section 5.5). The high memory efficiency of warping also enables high-resolution indexing, leading to improved accuracy.
@@ 4 Method
Figure 4: Visualizations on Spring, KITTI, and Sintel public benchmarks (from left to right).
@@
We develop two ways to adapt large-scale pre-trained models. Adaptation 1 (a1) is our initial design specific to DAv2. We freeze the entire DAv2, incorporate features from its DPT head, and further refine these features using a ResNet18. Adaptation 2 (a2) is our improved design that works better and supports more backbones, where we only freeze the ViT/CNN backbones. We make the DPT head trainable and side-tune the features with 3 ResNet blocks. For completeness, we report the results of both.
@@
Similar to existing iterative methods (Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022), our recurrent update module $R$ iteratively predicts the residual flow updates. At step $t$, we concatenate $F(I_{1})$ (feature of frame 1), $\texttt{Warp}(f_{\text{cur}})$ (warped feature of frame 2, see Section 3.3), and the current hidden state $\text{Hidden}_{t}\in\mathbb{R}^{h\times w\times d}$ as input. We use a slightly modified DPT (Ranftl et al., 2021) as the architecture of the module.
@@
We adopt the Mixture-of-Laplace (MoL) loss used in SEA-RAFT (Wang et al., 2024). At step $t$, the hidden state $\text{Hidden}_{t}$ is used to predict the MoL parameters $M\in\mathbb{R}^{h\times w\times 6}$. They are upsampled to the original image resolution through convex upsampling (Teed & Deng, 2020; Wang et al., 2024).
@@
We replace cost volumes, which are standard in existing iterative methods (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024), with high-resolution warping, which is more memory-efficient. In addition, we have removed the context encoder (Sun et al., 2018a), another flow-specific design standard in existing iterative methods.
A direct benefit from our simplified design is that we can load pre-trained weights for standard architectures such as ViT (Dosovitskiy et al., 2020), which can improve generalization as our experiments will show Section 5.5.
The simplicity of our meta-architecture also enables more apples-to-apples comparisons between direct methods and iterative methods. Existing direct methods (Weinzaepfel et al., 2022; 2023; Saxena et al., 2024) share an input format similar to that of the first iteration of WAFT, making them more directly compatible. We will empirically show the effectiveness and necessity of iterative indexing within our meta-architecture in Section 5.5.